OpenStudy (anonymous):

Simplify: (csc x)(sin ^2)(pi/2 -x)(tan x)

OpenStudy (anonymous):

I keep trying to apply the trig identities but it just gets me nowhere.

OpenStudy (anonymous):

:) hi again.

OpenStudy (whpalmer4):

is that \[\csc x*\sin^2(\pi/2-x)*\tan x\]?

OpenStudy (anonymous):

yup.

OpenStudy (whpalmer4):

What does the graph of \[y=\sin(\pi/2-x)\] look like compared to the graph of \[y=\sin(-x)\]

OpenStudy (anonymous):

I dont know lemme graph it...

OpenStudy (anonymous):

sin(-x) is shifted to the left

OpenStudy (anonymous):

No i guess it wouldbe the first one thats been shifted.

OpenStudy (anonymous):

:( Am I missinig something obvious again?

OpenStudy (whpalmer4):

Does it look like the graph of any other function?

OpenStudy (anonymous):

y=sin x

OpenStudy (whpalmer4):

\[\sin(\pi/2-x) = \cos(x)\] x pi/2-x sin pi/2-x cos x 0 pi/2 1 1 pi/2 0 0 0 pi -pi/2 -1 -1 etc.

OpenStudy (whpalmer4):

If we make that substitution, what do we have?

OpenStudy (anonymous):

but its a sin squared. not just a plain ol' sin.

OpenStudy (whpalmer4):

that's fine. if sin(pi/2-x) = cos(x) then sin^2(pi/2-x) = cos^x

OpenStudy (whpalmer4):

it's not a different function when squared, it's the same function, squared.

OpenStudy (anonymous):

okay... so then it'd be csc x cos^2 x tan x

OpenStudy (whpalmer4):

\[\sin^2x = (\sin x)*(\sin x)\]

OpenStudy (whpalmer4):

oops, didn't get the exponent in for cos^2x a few posts ago

OpenStudy (anonymous):

I figured it was supposed to be there :)

OpenStudy (anonymous):

Ugh i hate these problems I cant even work through them! I just sit here staring at it.

OpenStudy (whpalmer4):

What was the definition of csc again?

OpenStudy (anonymous):

1/sin x

OpenStudy (anonymous):

Okay I get to about \[\frac{ (\cos ^{2} x)(\tan x) }{ \sin x }\]

OpenStudy (anonymous):

Then I'm stuck and I have no idea if thats even right.

OpenStudy (whpalmer4):

Good. Do you know the definition of tan x?

OpenStudy (anonymous):

tan=sin/cos

OpenStudy (whpalmer4):

Okay, why don't you plug that in as well

OpenStudy (anonymous):

is it seriously that freakin simple? I'm going to shoot myself.

OpenStudy (whpalmer4):

the real stumbling block is realizing that sin(pi/2-x) = cos(x)

OpenStudy (anonymous):

\[\frac{ \cos ^{2} x}{ \sin x }\times \frac{ \sin x }{ \cos x }\]

OpenStudy (anonymous):

so then it =cos x

OpenStudy (whpalmer4):

yep, cos x is the final answer. Forehead, look out, incoming palm! :-)

OpenStudy (anonymous):

lol no joke...Do you wanna help with the next one too? Maybe if we stick together we'll just get super fast.

OpenStudy (whpalmer4):

Give it a shot first, then post it, there's someone else pleading for help...

OpenStudy (anonymous):

lol okay. :) us poor retarded people.